Explain the temperature of a junction of combination of two rods with derivation.

Grade:11

1 Answers

Vikas TU
7 years ago
U need to find the Resistance first for both rods.
from R=pl/A
then use,
H1 = H2
that is,
(T1-T)R1 = (T2-T)/R2